# Polarized QED Cascades over Pulsar Polar Caps

### v1

@@ Abstract
The formation of $e^{\pm}$ plasmas within pulsar magnetospheres through quantum electrodynamics (QED) cascades in vacuum gaps is widely acknowledged. This paper aims to investigate the effect of photon polarization during the QED cascade occurring over the polar cap of a pulsar. We employ a Monte Carlo-based QED algorithm that accurately accounts for both spin and polarization effects during photon emission and pair production in both single-particle and particle-in-cell (PIC) simulations. Our findings reveal distinctive properties in the photon polarization of curvature radiation (CR) and synchrotron radiation (SR). CR photons exhibit high linear polarization parallel to the plane of the curved magnetic field lines, whereas SR photons, on average, demonstrate weak polarization. As the QED cascade progresses, SR photons gradually dominate over CR photons, thus reducing the average degree of photon polarization. Additionally, our study highlights an intriguing observation: the polarization of CR photons enhances $e^{\pm}$ pair production by approximately 5%, in contrast to the inhibition observed in laser-plasma interactions. Our self-consistent QED PIC simulations in the corotating frame reproduce the essential results obtained from single-particle simulations.
@@
pulsar magnetosphere – QED cascade – electron-positron pair – gamma photon – spin and polarization effects
@@ 1 Introduction
Figure 1: Schematic representation of the polarized QED cascade over a pulsar polar cap. High-energy $e^{\pm}$ particles, traversing the curved magnetic field, emit $\gamma_{\parallel}$ photons through CR, which subsequently convert into $e^{\pm}$ pairs. The resulting $e^{\pm}$ pairs can emit both $\gamma_{\parallel}$ and $\gamma_{\perp}$ photons via SR. The symbols $\gamma_{\parallel}$ and $\gamma_{\perp}$ denote photons polarized parallel and perpendicular to the plane of the magnetic field lines, respectively.
The polarized QED cascade process over a pulsar polar cap is illustrated in Fig. 1. High-energy $e^{\pm}$, accelerated by the electric field, are confined to move along the curved magnetic field lines with minimal pitch angles, emitting $\gamma$ photons through curvature radiation (CR). These emitted $\gamma$ photons travel straight, leading to an increase in their propagation angles relative to the magnetic field lines until they decay into $e^{\pm}$ pairs. The created $e^{\pm}$ pairs, inheriting large pitch angles from their parent $\gamma$ photons, undergo synchrotron radiation (SR) emission while simultaneously experiencing strong radiation reaction. Subsequently, SR photons with sufficient energy also decay into $e^{\pm}$ pairs. The unscreened parallel electric field $E_{\parallel}$ accelerates the created $e^{\pm}$ pairs, reducing their pitch angles and potentially bringing them into the CR regime. This cycle of photon emission and pair production sustains the QED cascade. Ultimately, the generated $e^{\pm}$ plasma becomes sufficiently dense to screen the parallel electric field $E_{\parallel}$, thus creating the pulsar magnetosphere and approaching a force-free environment.
Magnetohydrodynamics simulations (Spitkovsky, 2006) and particle-in-cell (PIC) simulations based on heuristic QED models (Philippov & Spitkovsky, 2014; Chen & Beloborodov, 2014; Belyaev, 2015; Cerutti et al., 2015; Kalapotharakos et al., 2018; Timokhin & Harding, 2019; Cruz et al., 2022) have been conducted to replicate force-free pulsar magnetosphere features (Contopoulos et al., 1999). However, to accurately capture local dynamics like the QED cascade, first-principles QED-PIC simulations that incorporate precise rates of photon emission and pair production are essential. Previous 1D QED-PIC simulations in the corotating frame suggest that the QED cascade over the pulsar polar cap is dynamic, contrary to previous assumptions of stationarity (Timokhin, 2010; Timokhin & Arons, 2013). This nonstationary creation of $e^{\pm}$ plasmas may hold significance in explaining the origins of pulsar radio emission (Philippov et al., 2020; Cruz et al., 2021).
@@
In this paper, we employ a semiclassical Monte Carlo-based single-particle code and a first-principles QED PIC code that resolves $e^{\pm}$ spin and $\gamma$ photon polarization to simulate the polarized QED cascade over a pulsar polar cap. Our simulation results show that emitted CR photons exhibit significant linear polarization parallel to the plane of the curved magnetic field line. This polarization arises from the $e^{\pm}$ acceleration, which determines the polarization of the emitted photons, being predominantly directed along the magnetic field plane. In contrast, the fast gyration motion of $e^{\pm}$ with nearly uniformly rotated acceleration directions results in a weak average polarization of SR photons. As CR photons, chiefly emitted by the primary $e^{\pm}$, are more energetic than SR photons, the average photon polarization increases with higher photon energy. The high degree of polarization of CR photons leads to a higher pair production rate compared to the unpolarized one, resulting in approximately 5% more $e^{\pm}$ pairs in simulations accounting for photon polarization. Both “shower-type” (S-type), i.e. without parallel electric field $E_{\parallel}$, and self-sustained “avalanche-type” (A-type), i.e. with $E_{\parallel}$, QED cascades exhibit similar features, with A-type cascades having a larger proportion of SR photons due to accelerated secondary $e^{\pm}$ pairs reaching higher energies. Fully 3D single-particle simulation results are confirmed by 1D QED-PIC simulations in the corotating frame.
@@
The remaining sections of this paper are organized as follows. Section 2 provides an overview of the strong-field QED theory and simulation method utilized. In Section 3, we conduct a detailed analysis of photon polarization in three distinct radiation regimes using single-particle simulations. Section 4 delves into the comprehensive cascade of photon emission and pair production, employing both single-particle simulations and QED-PIC simulations. Specifically, we study the time evolution of photon emission and analyze the impact of polarization on pair production. Finally, Section 5 offers a concise summary of our key findings.
@@
Figure 2: The normalized rates of photon emission, $W_{\gamma}/\overline{W}_{\gamma}$ (three left-most panels), and pair production, $W_{\pm}/\overline{W}_{\pm}$ (right-most panel), are calculated using Eqs.(1) and(2), where $\overline{W}_{\gamma}$ and $\overline{W}_{\pm}$ are the corresponding unpolarized rates. In the three left panels, the orange line illustrates the average photon polarization, $\overline{\xi}_{3}$, as a function of the photon energy ratio, $\varepsilon_{\gamma}/\varepsilon_{e}$. The QED parameters in these panels are set as $\chi_{e}=3\times 10^{-4}$, $\chi_{e}=0.2$, and $\chi_{e}=20$, while the right-most panel displays results with $\chi_{\gamma}=0.2$. In the right-most panel, observe the symmetry with respect to the $\varepsilon_{\gamma}/\varepsilon_{e}=0.5$ axis, and the increased probability of particle creation with $\xi^{\prime}_{3}<0$.
@@ 2 Simulation method
In accordance with the strong-field QED theory of polarization-dependent pair production (as outlined below), the velocity of primary particles responsible for CR emission is nearly aligned with the magnetic field lines. Consequently, the spin of the created $e^{\pm}$ pairs is essentially perpendicular to the direction of the magnetic field [^1] This stands in stark contrast to scenarios such as storage rings or laser-plasma interactions, where the velocity of primary particles is typically transverse to the external field, and the spin of the created $e^{\pm}$ pairs is predominantly aligned with the external field.. Hence, the fast spin precession of $e^{\pm}$ within the superstrong magnetic field of the pulsar results into a rapid depolarization. Indeed, in our fully spin- and polarized-resolved simulations, we observe no significant net $e^{\pm}$ spin polarization. Consequently, our focus shifts to the analysis of photon polarization. With resolved photon polarization, the semiclassical rates for photon emission and pair production in the LCFA can be written as (Baier et al., 1998)
$\frac{dW_{\gamma}}{d\varepsilon_{\gamma}}=$ $\frac{C_{0}\Delta t}{\varepsilon_{e}^{2}}\left[\frac{\varepsilon_{e}^{2}+\varepsilon_{e}^{\prime 2}}{\varepsilon_{e}\varepsilon_{e}^{\prime}}K_{\frac{2}{3}}(y_{1})-{\rm Int}K_{\frac{1}{3}}(y_{1})+\xi_{3}K_{\frac{2}{3}}(y_{1})\right],$ (1)
$\frac{dW_{\pm}}{d\varepsilon_{\pm}}$ $=\frac{C_{0}\Delta t}{\varepsilon_{\gamma}^{2}}\left[\frac{\varepsilon_{+}^{2}+\varepsilon_{-}^{2}}{\varepsilon_{+}\varepsilon_{-}}K_{\frac{2}{3}}(y_{2})+{\rm Int}K_{\frac{1}{3}}(y_{2})-\xi_{3}^{\prime}K_{\frac{2}{3}}(y_{2})\right],$ (2)
@@
where $\phi$ is the angle between the two transverse acceleration directions of $\bm{e}_{1}$ and $\bm{e}_{1}^{\prime}$. If $\bm{e}_{1}$ and $\bm{e}_{1}^{\prime}$ are collinear, then $\phi=0$ or $\pi$, and $\xi_{3}^{\prime}=\xi_{3}$; otherwise, $\xi_{3}^{\prime}=-\xi_{3}$ if $\bm{e}_{1}$ is perpendicular to $\bm{e}_{1}^{\prime}$. We will see that the latter case holds for CR photons.
The preferred polarization of emitted photons can be determined from Eq. (1). The polarized photon emission rates $W_{\gamma}$ normalized to the unpolarized ones $\overline{W}_{\gamma}$ are displayed in the three left panels of Fig. 2. Three QED parameters $\chi_{e}=3\times 10^{-4}$, $\chi_{e}=0.2$, and $\chi_{e}=20$ are shown, which correspond to CR, SR, and super-SR (defined in the next section), respectively. For all three regimes, the normalized photon emission rate $W_{\gamma}/\overline{W}\gamma$ increases with the photon polarization $\xi_{3}$ for fixed photon energy. This indicates that the emitted photons with $\xi_{3}>0$ dominate, being mainly linearly polarized along the $\bm{e}_{1}$ axis, i.e., the transverse acceleration axis. Note that since $\chi_{e}$ of CR is much smaller than 1, high-energy photons are still much less energetic than their parent $e^{\pm}$. In each panel, the average photon polarization $\overline{\xi}_{3}$ as a function of the photon energy ratio $\varepsilon_{\gamma}/\varepsilon_{e}$ is presented by the orange line. For CR, $\overline{\xi}_{3}$ grows from 0.5 at $\varepsilon_{\gamma}/\varepsilon_{e}\lesssim 10^{-6}$ to nearly 1 at $\varepsilon_{\gamma}/\varepsilon_{e}\approx 10^{-2}$. Thus, the highest energy CR photons are highly linearly polarized. For SR, where the emitted photons can take away most of the energy of their parent $e^{\pm}$, $\overline{\xi}_{3}$ also increases from roughly 0.5, reaching a maximum of about 0.7 at $\varepsilon_{\gamma}/\varepsilon_{e}\approx 0.3$ in the case of $\chi_{e}=0.2$, and then drops to 0 when $\varepsilon_{\gamma}/\varepsilon_{e}$ approaches 1. This is because the photon polarization is highly sensitive to the spin polarization of their parent $e^{\pm}$ when the emitted photon energy is close to the energy of the parent $e^{\pm}$. Therefore, the opposite result to CR photons is concluded: the highest energy SR photons are weakly polarized. For super-SR with a large $\chi_{e}$, e.g., $\chi_{e}=20$, the photon polarization peak at middle energies nearly disappears, leading to $\overline{\xi}_{3}$ almost monotonically decreasing from 0.5 to 0 with the increase of $\varepsilon_{\gamma}/\varepsilon_{e}$. Note that, as it will be clear below, super-SR often occurs for large values of $\chi_{e}$, but does not coincide with $\chi_{e}\gg 1$.
@@
Note that the photon polarizations $\xi_{3}$ and $\xi_{3}^{\prime}$ in Eqs. (1) and (2) are defined with respect to different axes, i.e., $\bm{e}_{1}$ and $\bm{e}_{1}^{\prime}$, respectively. For convenience, hereafter, we use the same basis $({\bm{e}}_{y},{\bm{e}}_{z},{\bm{e}}_{x})$ for $\gamma$ photons propagating along the assumed magnetic field direction ${\bm{e}}_{x}$. Therefore, the positive (negative) value of $\xi_{3}$ represents linear photon polarization mainly along the $\bm{e}_{y}$ ($\bm{e}_{z}$) direction.
@@ 3 Photon polarization
The polarization characteristics of CR and SR, often referred to as synchro-curvature radiation (Cheng & Zhang, 1996; Viganò et al., 2015), can be analyzed in a unified manner by considering an electron beam with an initial energy $\varepsilon_{0}$ and pitch angle $\theta_{0}$ injected into a purely curved magnetic field. In this case, we assume a simple field configuration where the magnetic field is predominantly directed along the $+\bm{e}_{x}$ direction and slightly curved towards the $+\bm{e}_{y}$ direction, with a constant magnitude of $B_{0}=4\times 10^{11}$ G. To reduce computational costs, we adopt a constant curvature radius of $\rho_{0}=10$ km, which is smaller than the typical values (100–1000 km) but comparable to that of offset polar caps (Harding & Muslimov, 2011). We set the initial pitch angle of the electron beam towards the $+\bm{e}_{y}$ direction, specifically $\theta_{y}|_{t=0}=\theta_{0}$ and $\theta_{z}|_{t=0}=0$, where $\theta_{y,z}=\arctan(p_{y,z}/|p_{x}|)$.
To confine the charged $e^{\pm}$ moving along the curved magnetic field, the net Lorentz force must act as a centripetal force, pointing from the $e^{\pm}$ position to the curvature center of the magnetic field line. The Lorentz force acting on $e^{\pm}$ is determined by its pitch angle. Therefore, the pitch angle $\theta$ of $e^{\pm}$ must be asymmetric about the magnetic field line during its propagation, a phenomenon known as drift velocity (Alfvén, 1963). The center of the pitch angle in the $\theta_{y}$-$\theta_{z}$ space will drift by $\theta_{d}=({\rm max}|\theta_{z}|-{\rm max}|\theta_{y}|)$, which gives the the drift angle (Alfvén, 1963)
@@
Photon emission attributed to CR occurs only for $e^{\pm}$ at small pitch angles $\theta\lesssim 2\theta_{\rm d}$, while SR dominates if $\theta\gg 2\theta_{\rm d}$. The defined boundary between these two mechanisms is essentially the same as that based on radiation power (Cheng & Zhang, 1996; Viganò et al., 2015). This is because the radiation power is directly related to the QED parameter $\chi_{e}\approx\gamma_{0}\theta B_{0}/B_{\rm cr}$, which is proportional to $\theta$ at $\theta\ll 1$, where $B_{\rm cr}=m_{e}^{2}c^{2}/e\hbar\approx 4.41\times 10^{13}$ G is the QED critical strength of the magnetic field.
@@
$\begin{cases}{\rm CR}:\theta<2\theta_{d};\\ {\rm SR}:\theta>2\theta_{d}~{}{\rm and}~{}\tau_{\gamma}>\tau_{g};\\ \operatorname{Super-SR}:\tau_{\gamma}<\tau_{g}.\end{cases}$ (5)
@@
(i) CR. With an initial pitch angle $\theta_{0}=10^{-8}$ and energy $\varepsilon_{0}=3$ TeV ($2\theta_{d}\approx 5.0\times 10^{-8}$), the photon emission resides in the CR regime. Due to the magnetic field’s curvature, the pitch angle is not symmetric about the field line; instead, its center shifts along the $-\bm{e}_{z}$ direction by $\theta_{\rm d}$, as shown in the top-left panel of Fig. 4. Since the QED parameter $\chi_{e}$, which determines the photon emission rate, is proportional to the pitch angle, most photons are emitted at $\theta_{z}<0$. According to the polarized photon emission rate, photons are predominantly linearly polarized along the transverse acceleration direction of $e^{\pm}$. Consequently, on average CR photons are linearly polarized along the $\bm{e}_{y}$ direction, resulting in an average photon polarization $\overline{\xi}_{3}>0$ in the chosen basis $({\bm{e}}_{y},{\bm{e}}_{z},{\bm{e}}_{x})$. Consistent with the theoretical expectation, $\overline{\xi}_{3}$ monotonically increases with the photon energy.
@@
(ii) SR. When the initial pitch angle is $\theta_{0}=10^{-6}$ and the energy is $\varepsilon_{0}=0.1$ TeV ($2\theta_{d}\approx 1.7\times 10^{-9}$), and simultaneously the gyration period $\tau_{g}$ is smaller than the photon emission time $\tau_{\gamma}$, the photon emission transitions into the SR regime. A notable difference from CR is that the pitch angle of SR is essentially symmetric about the magnetic field line in the $\theta_{y}$-$\theta_{z}$ space, as shown in the middle-left panel of Fig. 4. Consequently, photons are nearly uniformly emitted during the $e^{\pm}$ gyration motion. Although photons remain linearly polarized along the azimuthal direction, i.e., the acceleration direction, the overall polarization at different pitch angles nearly cancels out. The curvature effect of the magnetic field induces a weak inhomogeneity of the photon emission in the $\bm{e}_{z}$ direction, leading to a slight dominance of $\xi_{3}>0$ photons. The average polarization $\overline{\xi}_{3}$ for all photons is approximately 8%, considerably smaller than the CR polarization degree and almost insensitive to the photon energy. Thus, for computational convenience, we treat SR photons as weakly polarized and directly set the polarization of all SR photons to a constant value of 8% in our subsequent simulations, which allows us to use a numerical time-step larger than the gyration period.

### v2

@@ Abstract
The formation of $e^{\pm}$ plasmas within pulsar magnetospheres through quantum electrodynamics (QED) cascades in vacuum gaps is widely acknowledged. This paper aims to investigate the effect of photon polarization during the QED cascade occurring over the polar cap of a pulsar. We employ a Monte Carlo-based QED algorithm that accurately accounts for both spin and polarization effects during photon emission and pair production in both single-particle and particle-in-cell (PIC) simulations. Our findings reveal distinctive properties in the photon polarization of curvature radiation (CR) and synchrotron radiation (SR). CR photons exhibit high linear polarization parallel to the plane of the curved magnetic field lines, whereas SR photons, on average, demonstrate weak polarization. As the QED cascade progresses, SR photons gradually dominate over CR photons, thus reducing the average degree of photon polarization. Additionally, our study highlights an intriguing observation: the polarization of CR photons enhances $e^{\pm}$ pair production by approximately 5%, in contrast to the inhibition observed in laser-plasma interactions. Our self-consistent QED-PIC simulations in the corotating frame reproduce the essential results obtained from single-particle simulations.
@@
plasmas – polarization – radiation: dynamics – pulsars: general
@@ 1 Introduction
Figure 1: Schematic representation of the polarized QED cascade over a pulsar polar cap. High-energy $e^{\pm}$ particles, drifting in the curved magnetic field, emit $\gamma_{\parallel}$ photons through CR, which subsequently convert into $e^{\pm}$ pairs. The resulting $e^{\pm}$ pairs can emit both $\gamma_{\parallel}$ and $\gamma_{\perp}$ photons via SR. The symbols $\gamma_{\parallel}$ and $\gamma_{\perp}$ denote photons polarized parallel and perpendicular to the plane of the magnetic field lines, respectively.
The polarized QED cascade process over a pulsar polar cap is illustrated in Fig. 1. High-energy $e^{\pm}$, accelerated by the electric field, are confined to move along the curved magnetic field lines with minimal pitch angles, emitting $\gamma$ photons through curvature radiation (CR). These emitted $\gamma$ photons travel straight, leading to an increase in their propagation angles relative to the magnetic field lines until they decay into $e^{\pm}$ pairs. The created $e^{\pm}$ pairs, inheriting large pitch angles from their parent $\gamma$ photons, undergo synchrotron radiation (SR) while simultaneously experiencing strong radiation reaction. Subsequently, SR photons with sufficient energy also decay into $e^{\pm}$ pairs. The unscreened parallel electric field $E_{\parallel}$ accelerates the created $e^{\pm}$ pairs, reducing their pitch angles and potentially bringing them into the CR regime. This cycle of photon emission and pair production sustains the QED cascade. Ultimately, the generated $e^{\pm}$ plasma becomes sufficiently dense to screen the parallel electric field $E_{\parallel}$, thus creating the pulsar magnetosphere and approaching a force-free environment.
Magnetohydrodynamics simulations (Spitkovsky, 2006) and particle-in-cell (PIC) simulations based on heuristic QED models (Philippov & Spitkovsky, 2014; Chen & Beloborodov, 2014; Belyaev, 2015; Cerutti et al., 2015; Kalapotharakos et al., 2018; Timokhin & Harding, 2019; Cruz et al., 2022) have been conducted to replicate force-free pulsar magnetosphere features (Contopoulos et al., 1999). However, to accurately capture local dynamics like the QED cascade, first-principles QED-PIC simulations that incorporate precise rates of photon emission and pair production are essential. Previous 1D QED-PIC simulations in the corotating frame suggest that the QED cascade over the pulsar polar cap is dynamic (Timokhin, 2010; Timokhin & Arons, 2013), contrary to previous assumptions of stationarity. This nonstationary creation of $e^{\pm}$ plasmas may hold significance in explaining the origins of pulsar radio emission (Philippov et al., 2020; Cruz et al., 2021; Bransgrove et al., 2023).
@@
In this paper, we employ a semiclassical Monte Carlo-based single-particle code and a first-principles QED PIC code that resolves $e^{\pm}$ spin and $\gamma$ photon polarization to simulate the polarized QED cascade over a pulsar polar cap. Our simulation results show that emitted CR photons exhibit significant linear polarization parallel to the plane of the curved magnetic field line. This polarization arises from the $e^{\pm}$ acceleration, which determines the polarization of the emitted photons, being predominantly directed along the magnetic field plane. In contrast, the fast gyration motion of $e^{\pm}$ with nearly uniformly rotated acceleration directions results in a weak average polarization of SR photons. As CR photons, chiefly emitted by the primary $e^{\pm}$, are more energetic than SR photons, the average photon polarization increases with higher photon energy. The high degree of polarization of CR photons leads to a higher pair-production rate compared to the unpolarized one, resulting in approximately 5% more $e^{\pm}$ pairs in simulations accounting for photon polarization. Both “shower-type” (S-type), i.e. without parallel electric field $E_{\parallel}$, and self-sustained “avalanche-type” (A-type), i.e. with $E_{\parallel}$, QED cascades exhibit similar features, with A-type cascades having a larger proportion of SR photons due to accelerated secondary $e^{\pm}$ pairs reaching higher energies. Fully 3D single-particle simulation results are confirmed by 1D QED-PIC simulations in the corotating frame.
@@
The remaining sections of this paper are organized as follows. Section 2 provides an overview of the strong-field QED theory and simulation method utilized. In Section 3, we conduct a detailed analysis of photon polarization in three distinct radiation regimes using single-particle simulations. Section 4 delves into the comprehensive cascade of photon emission and pair production, employing both single-particle simulations and QED-PIC simulations. Specifically, we study the time evolution of photon emission and analyze the impact of photon polarization on pair production. Finally, Section 5 offers a concise summary of our key findings.
@@
Figure 2: The normalized rates of photon emission, $W_{\gamma}/\overline{W}_{\gamma}$ (three left-most panels), and pair production, $W_{\pm}/\overline{W}_{\pm}$ (right-most panel), are calculated using Eqs.(1) and (2), where $\overline{W}_{\gamma}$ and $\overline{W}_{\pm}$ are the corresponding unpolarized rates. In the three left panels, the orange line illustrates the average photon polarization, $\overline{\xi}_{3}$, as a function of the photon energy ratio, $\varepsilon_{\gamma}/\varepsilon_{e}$. The QED parameters in these panels are set as $\chi_{e}=3\times 10^{-4}$, $\chi_{e}=0.2$, and $\chi_{e}=20$, while the right-most panel displays results with $\chi_{\gamma}=0.2$. In the right-most panel, observe the symmetry with respect to the $\varepsilon_{\pm}/\varepsilon_{\gamma}=0.5$ axis, and the increased probability of particle creation with $\xi^{\prime}_{3}<0$.
@@ 2 Simulation method
In a pulsar polar cap the velocity of primary particles responsible for CR emission is nearly aligned with the magnetic field lines. Consequently, in accordance with the strong-field QED theory of polarization-dependent pair production (as outlined below), the spin of the created $e^{\pm}$ pairs is basically perpendicular to the direction of the magnetic field [^1] This stands in stark contrast to scenarios such as storage rings or laser-plasma interactions, where the velocity of primary particles is typically transverse to the external field, and the spin of the created $e^{\pm}$ pairs is predominantly aligned with the external field.. Hence, the fast spin precession of $e^{\pm}$ within the superstrong magnetic field of the pulsar results into a rapid depolarization. Indeed, in our fully spin- and polarized-resolved simulations, we observe no significant net $e^{\pm}$ spin polarization. Consequently, our focus shifts to the analysis of photon polarization. With averaged (initial state) and summed (final states) $e^{\pm}$ spin but resolved photon polarization, the semiclassical rates for photon emission and pair production in the LCFA can be written as (Baier et al., 1998)
$\frac{d^{2}W_{\gamma}}{d\varepsilon_{\gamma}dt}=$ $\frac{C_{0}}{\varepsilon_{e}^{2}}\left[\frac{\varepsilon_{e}^{2}+\varepsilon_{e}^{\prime 2}}{\varepsilon_{e}\varepsilon_{e}^{\prime}}K_{\frac{2}{3}}(y_{1})-{\rm Int}K_{\frac{1}{3}}(y_{1})+\xi_{3}K_{\frac{2}{3}}(y_{1})\right],$ (1)
$\frac{d^{2}W_{\pm}}{d\varepsilon_{\pm}dt}$ $=\frac{C_{0}}{\varepsilon_{\gamma}^{2}}\left[\frac{\varepsilon_{+}^{2}+\varepsilon_{-}^{2}}{\varepsilon_{+}\varepsilon_{-}}K_{\frac{2}{3}}(y_{2})+{\rm Int}K_{\frac{1}{3}}(y_{2})-\xi_{3}^{\prime}K_{\frac{2}{3}}(y_{2})\right],$ (2)
@@
where $\phi$ is the angle between the two transverse acceleration directions of $\bm{e}_{1}$ and $\bm{e}_{1}^{\prime}$. If $\bm{e}_{1}$ and $\bm{e}_{1}^{\prime}$ are collinear, then $\phi=0$ or $\pi$, and $\xi_{3}^{\prime}=\xi_{3}$; otherwise, $\xi_{3}^{\prime}=-\xi_{3}$ if $\bm{e}_{1}$ is perpendicular to $\bm{e}_{1}^{\prime}$. We will see that the latter case holds for the decay of CR photons.
The preferred polarization of emitted photons can be determined from Eq. (1). The polarized photon emission rates $W_{\gamma}$ normalized to the unpolarized ones $\overline{W}_{\gamma}$ are displayed in the three left panels of Fig. 2. Three QED parameters $\chi_{e}=3\times 10^{-4}$, $\chi_{e}=0.2$, and $\chi_{e}=20$ are shown, which correspond to CR, SR, and super-SR (defined in the next section), respectively. For all three regimes, the normalized photon emission rate $W_{\gamma}/\overline{W}_{\gamma}$ increases with the photon polarization $\xi_{3}$ for fixed photon energy. This indicates that the emitted photons with $\xi_{3}>0$ dominate, being mainly linearly polarized along the $\bm{e}_{1}$ axis, i.e., the transverse acceleration axis. Note that since $\chi_{e}$ of CR is much smaller than 1, high-energy photons are still much less energetic than their parent $e^{\pm}$. In each panel, the average photon polarization $\overline{\xi}_{3}$ as a function of the photon energy ratio $\varepsilon_{\gamma}/\varepsilon_{e}$ is presented by the orange line. For CR, $\overline{\xi}_{3}$ grows from 0.5 at $\varepsilon_{\gamma}/\varepsilon_{e}\lesssim 10^{-6}$ to nearly 1 at $\varepsilon_{\gamma}/\varepsilon_{e}\approx 10^{-2}$. Thus, the highest energy CR photons are highly linearly polarized. For SR, where the emitted photons can take away most of the energy of their parent $e^{\pm}$, $\overline{\xi}_{3}$ also increases from roughly 0.5, reaching a maximum of about 0.7 at $\varepsilon_{\gamma}/\varepsilon_{e}\approx 0.3$ in the case of $\chi_{e}=0.2$, and then drops to 0 when $\varepsilon_{\gamma}/\varepsilon_{e}$ approaches 1. This is because the photon polarization is highly sensitive to the spin polarization of their parent $e^{\pm}$ when the emitted photon energy is close to the energy of the parent $e^{\pm}$. Therefore, the opposite result to CR photons is concluded: the highest energy SR photons are weakly polarized. For super-SR with a large $\chi_{e}$, e.g., $\chi_{e}=20$, the photon polarization peak at middle energies nearly disappears, leading to $\overline{\xi}_{3}$ almost monotonically decreasing from 0.5 to 0 with the increase of $\varepsilon_{\gamma}/\varepsilon_{e}$. Note that, as it will be clear below, super-SR often occurs for large values of $\chi_{e}$, but does not coincide with $\chi_{e}\gg 1$.
@@
Note that the photon polarizations $\xi_{3}$ and $\xi_{3}^{\prime}$ in Eqs. (1) and (2) are defined with respect to different axes, i.e., $\bm{e}_{1}$ and $\bm{e}_{1}^{\prime}$, respectively. For convenience, hereafter, we use the same basis $({\bm{e}}_{y},{\bm{e}}_{z},{\bm{e}}_{x})$ for $\gamma$ photons propagating along the assumed magnetic field direction ${\bm{e}}_{x}$. In the following, with this choice the basis, the positive (negative) value of $\xi_{3}$ represents linear photon polarization mainly along the $\bm{e}_{y}$ ($\bm{e}_{z}$) direction.
@@ 3 Photon polarization
The polarization characteristics of CR and SR, often referred to as synchro-curvature radiation (Cheng & Zhang, 1996; Viganò et al., 2015), can be analyzed in a unified manner by considering an electron beam with an initial energy $\varepsilon_{0}$ and pitch angle $\theta_{0}$ injected into a purely curved magnetic field. In this case, we assume a simple field configuration where the magnetic field is predominantly directed along the $+\bm{e}_{x}$ direction and slightly curved towards the $+\bm{e}_{y}$ direction, with a constant magnitude of $B_{0}=4\times 10^{11}$ G. To reduce computational costs, we adopt a constant curvature radius of $\rho_{0}=10$ km, which is smaller than the typical values (100–1000 km) but comparable to that of offset polar caps (Harding & Muslimov, 2011). We set the initial pitch angle of the electron beam towards the $+\bm{e}_{y}$ direction, specifically $\theta_{y}|_{t=0}=\theta_{0}$ and $\theta_{z}|_{t=0}=0$, where $\theta_{y,z}=\arctan(p_{y,z}/|p_{x}|)$, since photon emission occurs in the plane of the curved magnetic field lines ($x$-$y$ plane), preferentially.
To confine the charged $e^{\pm}$ moving along the curved magnetic field, the net Lorentz force must act as a centripetal force, pointing from the $e^{\pm}$ position to the curvature center of the magnetic field line. The Lorentz force acting on $e^{\pm}$ is determined by its pitch angle. Therefore, the pitch angle $\theta$ of $e^{\pm}$ must be asymmetric about the magnetic field line during its propagation, a phenomenon known as curvature drift velocity (Alfvén, 1963). The center of the pitch angle in the $\theta_{y}$-$\theta_{z}$ space will drift by $\theta_{d}=({\rm max}|\theta_{z}|-{\rm max}|\theta_{y}|)$, which gives the the drift angle (Alfvén, 1963)
@@
Photon emission attributed to CR occurs only for $e^{\pm}$ at small pitch angles $\theta\lesssim 2\theta_{\rm d}$, while SR dominates if $\theta\gg 2\theta_{\rm d}$. The defined boundary between these two mechanisms is essentially the same as that based on radiation power (Cheng & Zhang, 1996; Viganò et al., 2015), differing only by an unimportant factor of 2. This is because the radiation power is directly related to the QED parameter $\chi_{e}\approx\gamma_{0}\theta B_{0}/B_{\rm cr}$, which is proportional to $\theta$ at $\theta\ll 1$, where $B_{\rm cr}=m_{e}^{2}c^{2}/e\hbar\approx 4.41\times 10^{13}$ G is the QED critical strength of the magnetic field.
@@
$\begin{cases}{\rm CR}:\theta<2\theta_{d},\\ {\rm SR}:\theta>2\theta_{d}~{}{\rm and}~{}\tau_{\gamma}>\tau_{g},\\ \operatorname{super-SR}:\tau_{\gamma}<\tau_{g}.\end{cases}$ (5)
@@
(i) CR. With an initial pitch angle $\theta_{0}=10^{-8}$ and energy $\varepsilon_{0}=3$ TeV ($2\theta_{d}\approx 5.0\times 10^{-8}$), the photon emission resides in the CR regime. Due to the magnetic field’s curvature, the pitch angle is not symmetric about the field line; instead, its center shifts along the $-\bm{e}_{z}$ direction by $\theta_{\rm d}$, as shown in the top-left panel of Fig. 4. Since the QED parameter $\chi_{e}$, which determines the photon emission rate, is proportional to the pitch angle, most photons are emitted at $\theta_{z}<0$. According to the polarized photon emission rate, photons are predominantly linearly polarized along the net transverse acceleration direction of $e^{\pm}$. Consequently, on average CR photons are linearly polarized along the $\bm{e}_{y}$ direction, resulting in an average photon polarization $\overline{\xi}_{3}>0$ in the chosen basis $({\bm{e}}_{y},{\bm{e}}_{z},{\bm{e}}_{x})$. Consistent with the theoretical expectation, $\overline{\xi}_{3}$ monotonically increases with the photon energy.
@@
(ii) SR. When the initial pitch angle is $\theta_{0}=10^{-6}$ and the energy is $\varepsilon_{0}=0.1$ TeV ($2\theta_{d}\approx 1.7\times 10^{-9}$), and simultaneously the gyration period $\tau_{g}$ is smaller than the photon emission time $\tau_{\gamma}$, the photon emission transitions into the SR regime. A notable difference from CR is that the pitch angle of SR is essentially symmetric about the magnetic field line in the $\theta_{y}$-$\theta_{z}$ space, as shown in the middle-left panel of Fig. 4. Consequently, photons are nearly uniformly emitted during the $e^{\pm}$ gyration motion. Although photons remain linearly polarized along the acceleration direction, the overall polarization at different pitch angles nearly cancels out. The curvature effect of the magnetic field induces a weak inhomogeneity of the photon emission in the $\bm{e}_{z}$ direction, leading to a slight dominance of $\xi_{3}>0$ photons. The average polarization $\overline{\xi}_{3}$ for all photons is approximately 8%, considerably smaller than the CR polarization degree and almost insensitive to the photon energy. Thus, for computational convenience, we treat SR photons as weakly polarized and directly set the polarization of all SR photons to a constant value of 8% in our subsequent simulations, which allows us to use a numerical time-step larger than the gyration period.